OpenStudy (anonymous):

A Web music store offers two versions of a popular song.The size of the standard version is 2.2MB. The size of the high quality version is 4.2MB. Yesterday, there were 1250 downloads of the song, for a total size of 4290MB. How many downloads of the high quality version were there?

OpenStudy (nurali):

Let x= number of standard versions downloaded 1250-x= number of HQ versions downloaded 2.2x+4.2(1250-x)=4290 (in MB ) 2.2x+5250-4.2x=4290 2x=5250-4290 2x=960 x=480 There were 480 downloads of the HQ version

OpenStudy (anonymous):

Thanks!

OpenStudy (nurali):

My Pleasure.